### v1

# An Erdős problem on random subset sums in finite abelian groups

Abstract.
Let $f(N)$ denote the least integer $k$ such that, if $G$ is an abelian group of order $N$ and $A\subseteq G$ is a uniformly random $k$-element subset, then with probability at least $\tfrac{1}{2}$ the subset-sum set $\{\sum_{x\in S}x:S\subseteq A\}$ equals $G$. In 1965, Erdős and Rényi proved that for all $N$,
$f(N)\leq\log_{2}N+\mathopen{}\mathclose{{\left(\frac{1}{\log 2}+o(1)}}\right)\log\log N.$
Erdős later conjectured that this bound cannot be improved to $f(N)\leq\log_{2}N+o(\log\log N)$. In this paper we confirm this conjecture by showing that, for primes $p$,
$f(p)\geq\log_{2}p+\mathopen{}\mathclose{{\left(\frac{1}{2\log 2}+o(1)}}\right)\log\log p.$
This work is an outcome of human–AI collaboration: the original qualitative proof was generated autonomously by ChatGPT-5.2 Pro, while the quantitative refinement was developed by the authors.
Key words and phrases:
random subset sums, Erdős problems, abelian groups, second moment method
2020 Mathematics Subject Classification:
11B30, 60C05, 05D40
1. Introduction
The study of random subset sums in finite abelian groups dates back to the work of Erdős and Rényi [8] and of Erdős and Hall [7]. Let $f(N)$ denote the least integer $k$ such that the following holds: if $G$ is an abelian group of order $N$ and $A\subseteq G$ is a uniformly random $k$-element subset, then with probability at least $1/2$ every element of $G$ can be written as a sum of distinct elements of $A$. Erdős and Rényi [8] proved the universal upper bound
$f(N)\leq\log_{2}N+O(\log\log N),$ (1)
whereas Erdős and Hall [7] showed that this cannot, in general, be improved to
$f(N)\leq\log_{2}N+o(\log\log\log N).$
In 1973, Erdős [6, p. 127] conjectured that the $O(\log\log N)$ error term in (1) cannot be sharpened to $o(\log\log N)$. This conjecture is listed as Problem #543 on Bloom’s Erdős Problems website [4] in the following self-contained form.
Problem 1.1.
Define $f(N)$ to be the minimal $k$ such that the following holds: if $G$ is an abelian group of size $N$ and $A\subseteq G$ is a uniformly random $k$-element subset, then with probability at least $1/2$ every element of $G$ can be written in the form $\sum_{x\in S}x$ for some $S\subseteq A$. Is
$f(N)\leq\log_{2}N+o(\log\log N)\,?$
In this paper we confirm Erdős’ conjecture by answering Problem 1.1 in the negative. We restrict attention to prime orders $N=p$ (so $G$ has to be isomorphic to the cyclic group $\mathbb{F}_{p}$ of order $p$). This suffices to rule out the proposed bound $f(N)\leq\log_{2}N+o(\log\log N)$ for general $N$. The notion $\mathbb{F}_{p}$ usually stands for the field of $p$ elements, and here we only view it as an additive group. For any subset $A\subseteq\mathbb{F}_{p}$ we write
$\Sigma(A):=\Bigl\{\sum_{x\in S}x:\ S\subseteq A\Bigr\},$
where the empty sum is set to be 0. In this setting, equivalently, $f(p)$ denotes the least integer $k$ such that a uniformly random $k$-element subset $A\subseteq\mathbb{F}_{p}$ satisfies $\mathbb{P}(\Sigma(A)=\mathbb{F}_{p})\geq\tfrac{1}{2}$. Our main result is as follows.
Theorem 1.2.
Fix any constant $c$ with $0<c<\frac{1}{2\log 2}$. Let $p$ be prime and let $A\subseteq\mathbb{F}_{p}$ be a uniformly random $k$-element subset with $k=\lfloor\log_{2}p+c\log\log p\rfloor$. Then
$\mathbb{P}\bigl(\Sigma(A)=\mathbb{F}_{p}\bigr)\longrightarrow 0\qquad(p\to\infty).$
This provides a quantitative bound for primes, which answers Problem 1.1 in the negative.
Corollary 1.3.
For all sufficiently large primes $p$, it holds that
$f(p)\geq\log_{2}p+\mathopen{}\mathclose{{\left(\frac{1}{2\log 2}+o(1)}}\right)\log\log p.$
1.1. Declaration of AI usage
The interest of this work lies not only in the mathematical results presented but also in the process by which it was obtained, with large language models serving as a major research assistant. For transparency and reproducibility, we provide the AI-assisted workflow below. [^1] For a public record of the development, see the second author’s comments posted at 13:12 on 21 Jan 2026 and 11:37 on 23 Jan 2026 in the forum thread at https://www.erdosproblems.com/forum/thread/543.
  1. (i)
    In an initial chat, we asked ChatGPT-5.2 Pro to search for a counterexample to the Erdős problem stated in [4] (i.e., the inequality in Problem 1.1). The model returned a proposed counterexample together with an accompanying argument.
  2. (ii)
    In a separate chat, we asked ChatGPT-5.2 Pro to verify the proof from the first session. It reported a serious gap that invalidated the argument as written.
  3. (iii)
    We then asked the second chat session to repair the argument. It produced a patched version that indeed disproves the inequality in Problem 1.1. The AI-generated proof, with minimal edits for readability, is available at [1].
  4. (iv)
    Motivated by the upper bound due to Erdős and Rényi [8], we observed that for prime $p$ one can obtain a quantitative lower bound, rather than only the qualitative disproof produced by the AI. By strengthening several steps of the AI-generated argument, we derived the quantitative lower bound for $f(p)$ stated in Theorem 1.2.
We emphasize that step (iv), as well as the writing of the paper, was carried out by the human authors. Related papers on AI-assisted solutions to Erdős problems have appeared recently; see, for example, Sothanaphan [13], Barreto et al. [3], and Feng et al. [9]. In our view, this work serves as an illustrative case study of AI-assisted use of the probabilistic method and linear algebra in combinatorics.
1.2. Paper organization
The rest of the paper is organized as follows. Section 2 collects notation and several auxiliary lemmas. In Section 3 we prove Theorem 1.2 by first establishing the key Poisson-like estimate, Proposition 3.1, and then combining it with a second-moment argument. Section 4 contains several concluding remarks.
2. Preliminaries
2.1. Notation
Throughout, $\log$ denotes the natural logarithm and $\log_{2}x:=\log x/\log 2$. For a positive integer $k$, we define $k]:=\{1,\dots,k\}$. For $Y\in\mathbb{R}$ and $r\in\mathbb{N}$, let $(Y)_{r}:=Y(Y-1)\cdots(Y-r+1)$ denote the falling factorial.
We use Vinogradov’s asymptotic notation. For functions $f=f(n)$ and $g=g(n)$, we write $f=O(g)$, $g=\Omega(f)$, $f\ll g$, or $g\gg f$ to mean that there exists a constant $C>0$ such that $|f(n)|\leq Cg(n)$ for all sufficiently large $n$. We write $f\asymp g$ or $f=\Theta(g)$ to mean that $f\ll g$ and $g\ll f$, and we write $f=o(g)$ to mean that $f(n)/g(n)\to 0$ as $n\to\infty$.
2.2. Auxiliary lemmas
We now present several auxiliary lemmas that will be used repeatedly in the proof of the main theorem. We begin with two linear-algebraic observations for $\{0,1\}$-matrices. Let $\mathbb{Q}$ denote the field of rational numbers.
Lemma 2.1.
Fix a constant $\beta$ with $0<\beta<\tfrac{1}{2}$. Let $p$ be prime and let $V\in\{0,1\}^{r\times k}$ with $r\leq(\log p)^{\beta}$. Then for all sufficiently large primes $p$,
$\operatorname{rank}_{\mathbb{Q}}(V)=\operatorname{rank}_{\mathbb{F}_{p}}(V).$
Proof.
Let $s=\operatorname{rank}_{\mathbb{Q}}(V)$. Then there exists some $s\times s$ minor of $V$ with nonzero integer determinant $\Delta$. By Hadamard’s inequality,
$|\Delta|\leq s^{s/2}\leq r^{r/2}.$
Since $r\leq(\log p)^{\beta}$, we have
$\log\bigl(r^{r/2}\bigr)=\frac{r}{2}\log r\leq\frac{1}{2}(\log p)^{\beta}\log\log p=o(\log p),$
because $\beta<1$. Hence $r^{r/2}<p$ for all sufficiently large $p$, so $\Delta\not\equiv 0\pmod{p}$ and therefore $\operatorname{rank}_{\mathbb{F}_{p}}(V)\geq s$. The reverse inequality $\operatorname{rank}_{\mathbb{F}_{p}}(V)\leq\operatorname{rank}_{\mathbb{Q}}(V)$ holds for any integer matrix, since reducing modulo $p$ cannot increase rank. Thus $\operatorname{rank}_{\mathbb{F}_{p}}(V)=\operatorname{rank}_{\mathbb{Q}}(V)$. ∎
Lemma 2.2.
Let $V\in\{0,1\}^{r\times k}$ have distinct and nonzero rows. Let $W:=\mathrm{Col}(V)\subseteq\mathbb{Q}^{r}$ denote the column space of $V$ over $\mathbb{Q}$. Then its orthogonal complement $W^{\perp}$ contains no nonzero vector supported on at most two coordinates.
Proof.
Assume for contradiction that there exists some $y\in W^{\perp}\backslash\{0\}$ whose support has size at most $2$. We distinguish between the following two cases.
Case 1: $\mathrm{supp}(y)=\{i\}$. Then $y=ce_{i}$ for some $c\in\mathbb{Q}^{\times}$. Since $y\in W^{\perp}$, we have $\langle y,x\rangle=0$ for all $x\in W$, i.e. $c\,x_{i}=0$ for all $x\in W$, hence $x_{i}=0$ for all $x\in W$. In particular, taking $x$ to be any column of $V$, we obtain that the $i$th entry of every column is 0, so the $i$th row of $V$ is identically zero, contradicting the hypothesis.
Case 2: $\mathrm{supp}(y)=\{i,j\}$ with $i\neq j$. Then $y=\alpha e_{i}+\beta e_{j}$ for some $\alpha,\beta\in\mathbb{Q}^{\times}$. Again $y\in W^{\perp}$ implies
$\alpha x_{i}+\beta x_{j}=\langle y,x\rangle=0\qquad\text{for all }x\in W.$
Applying this to each column $v^{(t)}\in\mathbb{Q}^{r}$ of $V$ gives, for every $1\leq t\leq k$,
$\alpha\,V_{i,t}+\beta\,V_{j,t}=0.$
Equivalently, as vectors in $\mathbb{Q}^{k}$,
$\alpha\,V_{i,\ast}+\beta\,V_{j,\ast}=0,$
so $V_{i,\ast}=c\,V_{j,\ast}$ with $c:=-\beta/\alpha\in\mathbb{Q}^{\times}$, i.e. the $i$th and $j$th rows are proportional. Since $V_{j,\ast}\in\{0,1\}^{k}$ is nonzero, there exists an index $t$ with $V_{j,t}=1$, hence $c=V_{i,t}\in\{0,1\}$. As $c\neq 0$, we must have $c=1$, and therefore $V_{i,\ast}=V_{j,\ast}$, contradicting the assumption that the rows are distinct.
Hence no such $y$ exists in both cases, completing the proof of this lemma. ∎
Next, we provide a simple bound on the size of $W\cap\{0,1\}^{r}$ for proper subspaces $W\subseteq\mathbb{Q}^{r}$ under a mild non-degeneracy condition (i.e., the conclusion of Lemma 2.2).
Lemma 2.3.
Let $W\subseteq\mathbb{Q}^{r}$ be a $d$-dimensional subspace with $d<r$. If $W^{\perp}$ contains no nonzero vector supported on at most two coordinates, then
$|W\cap\{0,1\}^{r}|\leq\frac{3}{4}\,2^{d}.$
Proof.
Choose a coordinate set $J\subseteq[r]$ with $|J|=d$ such that the projection $\pi_{J}:W\to\mathbb{Q}^{J}$ is an isomorphism. Thus each $x\in W$ is uniquely determined by $u:=x|_{J}\in\mathbb{Q}^{d}$, and for each $i\notin J$ there exists a homogeneous linear form $L_{i}:\mathbb{Q}^{d}\to\mathbb{Q}$ with $x_{i}=L_{i}(u)$. Hence
$W\cap\{0,1\}^{r}\ \cong\ \Bigl\{u\in\{0,1\}^{d}:\ L_{i}(u)\in\{0,1\}\ \forall i\in[r]\setminus J\Bigr\}.$
We claim that there exists some $i_{0}\in[r]\setminus J$ such that $L_{i_{0}}$ has at least two nonzero coefficients. Indeed, suppose for contradiction that each $L_{i}$ depends on at most one coordinate (equivalently, its coefficient vector has support size at most $1$). Since each $L_{i}$ is a homogeneous linear form on $\mathbb{Q}^{d}$, for every $i\in[r]\setminus J$ we then have either
$L_{i}(u)\equiv 0,$
or else there exist $j(i)\in J$ and $c_{i}\in\mathbb{Q}^{\times}$ such that
$L_{i}(u)=c_{i}\,u_{j(i)}\qquad\text{for all }u\in\mathbb{Q}^{d}.$
In the first case, $x_{i}=L_{i}(x|_{J})\equiv 0$ for all $x\in W$, hence $e_{i}\in W^{\perp}$ (a nonzero vector of support $1$). In the second case, for all $x\in W$ we have $x_{i}=L_{i}(x|_{J})=c_{i}x_{j(i)}$, so $x_{i}-c_{i}x_{j(i)}\equiv 0$ on $W$ and therefore $e_{i}-c_{i}e_{j(i)}\in W^{\perp}$ (a nonzero vector of support $2$). In either case we obtain a nonzero vector in $W^{\perp}$ supported on at most two coordinates, contradicting the hypothesis. Thus there must exist $i_{0}\in[r]\setminus J$ for which $L_{i_{0}}$ has at least two nonzero coefficients.
Fix such $i_{0}$ and write $L_{i_{0}}(u)=\sum_{\ell=1}^{d}\alpha_{\ell}u_{\ell}$ with at least two $\alpha_{\ell}\neq 0$. Let $Z$ be uniform on $\{0,1\}^{d}$. Conditioning on all coordinates except two indices $s\neq t$ with $\alpha_{s},\alpha_{t}\neq 0$, the random variable $L_{i_{0}}(Z)$ becomes $C+aX+bY$ where $X,Y$ are independent $\mathrm{Bernoulli}(1/2)$ and $a,b\neq 0$. Thus $C+aX+bY$ takes four equiprobable values; in the degenerate case $a=\pm b$ it takes three values with masses $1/4,1/2,1/4$. In either case,
$\mathbb{P}\bigl(L_{i_{0}}(Z)\in\{0,1\}\bigr)\leq\frac{3}{4}.$
Therefore at most a $3/4$ fraction of $u\in\{0,1\}^{d}$ satisfy $L_{i_{0}}(u)\in\{0,1\}$, and imposing further constraints can only reduce the count. Hence $|W\cap\{0,1\}^{r}|\leq\frac{3}{4}\,2^{d}$. ∎
For the remainder of the paper, let $p$ be a sufficiently large prime and set
$k=\Bigl\lfloor\log_{2}p+c\log\log p\Bigr\rfloor,\qquad M:=2^{k}-1,\qquad\mbox{ and }\qquad\lambda:=\frac{M}{p},$
where $c\in\big(0,\frac{1}{2\log 2}\big)$ is arbitrary fixed. We will use the following estimate repeatedly.
Lemma 2.4.
Let $\alpha:=c\log 2\in(0,1/2)$. Then $\lambda=\Theta((\log p)^{\alpha})$. In particular,
$\lambda=o(\log p)\qquad\text{and}\qquad e^{\lambda}=p^{o(1)}.$
Proof.
Write $\theta_{p}:=\log_{2}p+c\log\log p-k\in[0,1)$. Then
$2^{k}=2^{\log_{2}p+c\log\log p-\theta_{p}}=p\,2^{c\log\log p}\,2^{-\theta_{p}}=p\,(\log p)^{c\log 2}\,2^{-\theta_{p}}=p\,(\log p)^{\alpha}\,2^{-\theta_{p}}.$
Thus $2^{k}/p$ lies between $(1/2)(\log p)^{\alpha}$ and $(\log p)^{\alpha}$. Since $M=2^{k}-1\sim 2^{k}$, we obtain $\lambda=M/p=\Theta((\log p)^{\alpha})$. The remaining claims follow immediately. ∎
We will also need the following crude but convenient estimate for falling factorials.
Lemma 2.5.
Fix a constant $\beta$ with $0<\beta<\tfrac{1}{2}$. Assume $m\in\{1,2\}$ and $r\leq(\log p)^{\beta}$. Then
$\frac{(mM)_{r}}{p^{r}}=m^{r}\,p^{o(1)}.$
Proof.
Write
$(mM)_{r}=\prod_{j=0}^{r-1}(mM-j)=(mM)^{r}\prod_{j=0}^{r-1}\Bigl(1-\frac{j}{mM}\Bigr).$
Since $0\leq j\leq r-1<(\log p)^{\beta}$, we have $j/(mM)=o(1)$ uniformly in $j$, and thus the fact $\log(1-x)=-x+O(x^{2})$ leads to
$\sum_{j=0}^{r-1}\log\Bigl(1-\frac{j}{mM}\Bigr)=-\frac{1}{mM}\sum_{j=0}^{r-1}j+O\Bigl(\frac{1}{(mM)^{2}}\sum_{j=0}^{r-1}j^{2}\Bigr).$
Using $\sum_{j<r}j=O(r^{2})$ and $\sum_{j<r}j^{2}=O(r^{3})$, we obtain
$\sum_{j=0}^{r-1}\log\Bigl(1-\frac{j}{mM}\Bigr)=O\Bigl(\frac{r^{2}}{M}\Bigr)+O\Bigl(\frac{r^{3}}{M^{2}}\Bigr)=o(1),$
because $r\leq(\log p)^{\beta}$ while $M=2^{k}-1\geq p^{1+o(1)}\to\infty$. Therefore
$(mM)_{r}=(mM)^{r}\exp(o(1))=(mM)^{r}\,(1+o(1)).$ (2)
Dividing by $p^{r}$ yields
$\frac{(mM)_{r}}{p^{r}}=m^{r}\Bigl(\frac{M}{p}\Bigr)^{r}(1+o(1)).$
By Lemma 2.4, we have $M/p=\lambda=\Theta((\log p)^{\alpha})$ and thus $\log(M/p)=\Theta(\log\log p)$. So
$\log\mathopen{}\mathclose{{\left(\Bigl(\frac{M}{p}\Bigr)^{r}}}\right)=r\log(M/p)=O((\log p)^{\beta}\log\log p)=o(\log p),$
implying that $(M/p)^{r}=p^{o(1)}$. Combining the above estimates gives
$\frac{(mM)_{r}}{p^{r}}=m^{r}\,p^{o(1)},$
as desired. ∎
3. Proof of Main Result
In this section we prove Theorem 1.2 in three stages. First, in Section 3.1, we reduce the problem to an independent (i.i.d.) model. The proof of the theorem itself (Section 3.2) is based on a second-moment analysis, which relies on a crucial Poisson-like estimate (Proposition 3.1). The proof of this proposition, which constitutes our main technical contribution, is given in Section 3.3 and proceeds by applying Bonferroni inequalities as well as estimating factorial moments of an associated Poisson-like random variable.
3.1. Reducing to an independent model
Fix a prime $p$ and an integer $k=O(\log p)$. We compare the following two distributions on subsets of $\mathbb{F}_{p}$. In the subset model, $A_{\mathrm{subset}}$ is chosen uniformly from $\binom{\mathbb{F}_{p}}{k}$, and for any event $F$ that depends only on the subset $A_{\mathrm{subset}}$, we write
$\mathbb{P}_{\mathrm{subset}}(F):=\mathbb{P}\bigl(F(A_{\mathrm{subset}})\bigr).$
In the independent (i.i.d.) model, let $a_{1},\dots,a_{k}$ be i.i.d. uniform on $\mathbb{F}_{p}$ (allowing repetitions) and set
$A_{\mathrm{iid}}:=\{a_{1},\dots,a_{k}\}\subseteq\mathbb{F}_{p},\qquad\mathbb{P}_{\mathrm{iid}}(F):=\mathbb{P}\bigl(F(A_{\mathrm{iid}})\bigr).$
Let $E$ be the event that $a_{1},\dots,a_{k}$ are pairwise distinct. Since $k=O(\log p)$, a union bound gives
$\mathbb{P}(E^{c})=\mathbb{P}(\exists\,i<j:\ a_{i}=a_{j})\leq\binom{k}{2}\frac{1}{p}=o(1)\qquad(p\to\infty).$ (3)
Conditioned on $E$, the unordered set $A_{\mathrm{iid}}$ is uniformly distributed over $\binom{\mathbb{F}_{p}}{k}$, so $A_{\mathrm{iid}}\mid E$ has the same distribution as $A_{\mathrm{subset}}$. By the law of total probability, for any event $F$ we have
$\bigl|\mathbb{P}(F\mid E)-\mathbb{P}(F)\bigr|\leq 2\mathbb{P}(E^{c}).$
Consequently, for every such event $F$, by (3) we have
$\bigl|\mathbb{P}_{\mathrm{subset}}(F)-\mathbb{P}_{\mathrm{iid}}(F)\bigr|=\bigl|\mathbb{P}(F(A_{\mathrm{iid}})\mid E)-\mathbb{P}(F(A_{\mathrm{iid}}))\bigr|\leq 2\,\mathbb{P}(E^{c})=o(1).$ (4)
Applying (4) to the event $F=\{\Sigma(A)=\mathbb{F}_{p}\}$ yields
$\mathbb{P}_{\mathrm{subset}}\bigl(\Sigma(A)=\mathbb{F}_{p}\bigr)\leq\mathbb{P}_{\mathrm{iid}}\bigl(\Sigma(A_{\mathrm{iid}})=\mathbb{F}_{p}\bigr)+o(1).$ (5)
Therefore, in order to prove Theorem 1.2 it suffices to show that
$\mathbb{P}_{\mathrm{iid}}\bigl(\Sigma(A_{\mathrm{iid}})=\mathbb{F}_{p}\bigr)\longrightarrow 0\qquad(p\to\infty).$
Henceforth all probabilities and expectations are taken in the i.i.d. model, and for convenience, we omit the subscript “iid” throughout the rest of the paper.
3.2. Proof of Theorem 1.2 via a Poisson-like estimate
To prove Theorem 1.2, we rephrase the event $F=\{\Sigma(A)=\mathbb{F}_{p}\}$ in terms of the number of “missed” elements in $\mathbb{F}_{p}$ as defined below. For each nonempty $S\subseteq[k]$, define the indexed subset sum
$\sigma(S):=\sum_{i\in S}a_{i}\in\mathbb{F}_{p}.$ (6)
[…]
$X_{x}:=\#\{\varnothing\neq S\subseteq[k]:\ \sigma(S)=x\}.$
Thus $X_{x}=0$ denotes the event that $x$ is missed by all nonempty indexed subset sums. Set
$U:=\#\{x\in\mathbb{F}_{p}^{\times}:\ X_{x}=0\}=\sum_{x\in\mathbb{F}_{p}^{\times}}\mathbf{1}_{\{X_{x}=0\}}.$
Since $0\in\Sigma(A)$ always (via the empty sum), we have $\Sigma(A)=\mathbb{F}_{p}\ \Longrightarrow\ U=0$, and consequently
$\mathbb{P}\bigl(\Sigma(A)=\mathbb{F}_{p}\bigr)\leq\mathbb{P}(U=0).$ (7)
The technical core of this section is as follows. For $B\subseteq\mathbb{F}_{p}^{\times}$, define $X_{B}:=\sum_{x\in B}X_{x}$. Intuitively speaking, one can view each $X_{x}$ as a Poisson random variable of mean $\lambda=\frac{M}{p}=\frac{2^{k}-1}{p}$, where each non-empty subset $S\subseteq[k]$ contributes the value $\sum_{i\in S}a_{i}$ uniformly in $\mathbb{F}_{p}$.
Proposition 3.1.
Fix $m\in\{1,2\}$. For any $B\subseteq\mathbb{F}_{p}^{\times}$ with $|B|=m$,
$\mathbb{P}\bigl(X_{B}=0\bigr)=e^{-m\lambda}\,(1+o(1))\qquad(p\to\infty),$
where the $o(1)$ is uniform over all such $B$.
Assuming Proposition 3.1, we now complete the proof of Theorem 1.2.
Proof of Theorem 1.2, assuming Proposition 3.1.
By Proposition 3.1 with $m=1$, for each fixed $x\in\mathbb{F}_{p}^{\times}$ we have
$\mathbb{P}(X_{x}=0)=e^{-\lambda}(1+o(1)).$
Recall that $U=\sum_{x\in\mathbb{F}_{p}^{\times}}\mathbf{1}_{\{X_{x}=0\}}$. By linearity of expectation, we have
$\mathbb{E}U=\sum_{x\in\mathbb{F}_{p}^{\times}}\mathbb{P}(X_{x}=0)=(p-1)e^{-\lambda}(1+o(1))=pe^{-\lambda}(1+o(1)).$ (8)
Similarly, for distinct $x\neq y$ in $\mathbb{F}_{p}^{\times}$, Proposition 3.1 (applied with $m=2$ and $B=\{x,y\}$) yields
$\mathbb{P}(X_{x}=0,\ X_{y}=0)=e^{-2\lambda}(1+o(1)),$
and hence we can derive that
$\mathbb{E}[U(U-1)]=\sum_{\begin{subarray}{c}x,y\in\mathbb{F}_{p}^{\times}\\ x\neq y\end{subarray}}\mathbb{P}(X_{x}=0,X_{y}=0)=(p-1)(p-2)e^{-2\lambda}(1+o(1))=p^{2}e^{-2\lambda}(1+o(1)).$ (9)
From (8)–(9) we obtain
$\frac{\mathbb{E}[U(U-1)]}{(\mathbb{E}U)^{2}}=\frac{p^{2}e^{-2\lambda}(1+o(1))}{p^{2}e^{-2\lambda}(1+o(1))}=1+o(1),$
and thus
$\frac{\mathrm{Var}(U)}{(\mathbb{E}U)^{2}}=\frac{\mathbb{E}U+\mathbb{E}[U(U-1)]-(\mathbb{E}U)^{2}}{(\mathbb{E}U)^{2}}=\frac{1}{\mathbb{E}U}+\mathopen{}\mathclose{{\left(\frac{\mathbb{E}[U(U-1)]}{(\mathbb{E}U)^{2}}-1}}\right)=\frac{1}{\mathbb{E}U}+o(1).$
By Lemma 2.4 we have $\lambda=o(\log p)$, so $\log(\mathbb{E}U)=\log p-\lambda+o(1)\to\infty$ and hence $\mathbb{E}U\to\infty$. Therefore $\mathrm{Var}(U)/(\mathbb{E}U)^{2}\to 0$, and Chebyshev’s inequality gives
$\mathbb{P}(U=0)\leq\mathbb{P}\bigl(|U-\mathbb{E}U|\geq\mathbb{E}U\bigr)\leq\frac{\mathrm{Var}(U)}{(\mathbb{E}U)^{2}}=o(1).$
Combining this with (7) and (5), we conclude
$\mathbb{P}_{\mathrm{subset}}\bigl(\Sigma(A)=\mathbb{F}_{p}\bigr)=o(1),$
which proves Theorem 1.2. ∎
The remainder of this section is devoted to the proof of Proposition 3.1.
3.3. Proof of Proposition 3.1
Fix $m\in\{1,2\}$ and $B\subseteq\mathbb{F}_{p}^{\times}$ with $|B|=m$. We estimate the factorial moments of $X_{B}$ by comparing them with those of a Poisson random variable of mean $m\lambda$, and then apply these bounds in the following Bonferroni-type inequalities (which arise from the inclusion-exclusion principle on factorial moments; see, e.g., [2, Lemma 2.1] and [5, Theorem 8.2.1(b)]). For every integer-valued random variable $Z\geq 0$, it holds that
$\mathbb{P}(Z=0)=\sum_{r=0}^{\infty}(-1)^{r}\frac{\mathbb{E}[(Z)_{r}]}{r!},$
and for every integer $R\geq 0$,
$\mathopen{}\mathclose{{\left|\mathbb{P}(Z=0)-\sum_{r=0}^{R}(-1)^{r}\frac{\mathbb{E}[(Z)_{r}]}{r!}}}\right|\leq\frac{\mathbb{E}[(Z)_{R+1}]}{(R+1)!}.$ (10)
Throughout the rest of this proof, choose an exponent $\beta$ with
$\alpha<\beta<\frac{1}{2},$
where $\alpha=c\log 2$ is the same as in Lemma 2.4, and set
$R:=\Bigl\lfloor(\log p)^{\beta}\Bigr\rfloor.$
Recall that $M=2^{k}-1$ and $\lambda=M/p$. By Lemma 2.4, $\lambda=o(R)$. It suffices to establish the following uniform estimate
$\mathbb{E}[(X_{B})_{r}]=(m\lambda)^{r}\,(1+o(1))\qquad(1\leq r\leq R).$ (11)
As we will show later, once this estimate is proved, substituting it into (10) with $Z=X_{B}$ completes the proof.
To proceed, we define some notations. Write $a=(a_{1},\dots,a_{k})\in\mathbb{F}_{p}^{k}$. For each nonempty $S\subseteq[k]$, let $v(S)\in\{0,1\}^{k}$ be its indicator vector. So the definition (6) becomes $\sigma(S)=a\cdot v(S)$. Given subsets $S_{1},\dots,S_{r}\subseteq[k]$, we write $V=V(S_{1},\dots,S_{r})\in\{0,1\}^{r\times k}$ for the incidence matrix with entries $V_{j,t}=\mathbf{1}_{\{t\in S_{j}\}}$.
The following lemma relates estimates of certain probabilities to corresponding ranks. Fix distinct nonempty subsets $S_{1},\dots,S_{r}$ and elements $b_{1},\dots,b_{r}\in\mathbb{F}_{p}$, and let $d:=\operatorname{rank}_{\mathbb{F}_{p}}(V)$.
Lemma 3.2.
With notation as above, the system $a\cdot v(S_{j})=b_{j}$ for $1\leq j\leq r$ is either inconsistent (probability 0), or consistent and then
$\mathbb{P}\bigl(\sigma(S_{j})=b_{j}\ \forall j\bigr)=p^{-d}.$
Proof.
Define the linear map $T:\mathbb{F}_{p}^{k}\to\mathbb{F}_{p}^{r}$ by $T(a)=(a\cdot v(S_{1}),\dots,a\cdot v(S_{r}))$. Then $\dim(\mathrm{Im}\,T)=\mathrm{rank}_{\mathbb{F}_{p}}(V)=d$, so $|\mathrm{Im}\,T|=p^{d}$ and every fiber has size $|\ker T|=p^{k-d}$. If $b:=(b_{1},\dots,b_{r})\notin\mathrm{Im}\,T$ then $\mathbb{P}(T(a)=b)=0$; otherwise $\mathbb{P}(T(a)=b)=p^{k-d}/p^{k}=p^{-d}$. ∎
Now we expand $(X_{B})_{r}$ by counting ordered $r$-tuples of distinct pairs $(S_{j},b_{j})$ with $\varnothing\neq S_{j}\subseteq[k]$ and $b_{j}\in B$:
$(X_{B})_{r}=\sum_{\begin{subarray}{c}(S_{1},b_{1}),\dots,(S_{r},b_{r})\\ \text{all distinct}\end{subarray}}\mathbf{1}_{\{\sigma(S_{1})=b_{1},\dots,\sigma(S_{r})=b_{r}\}}.$
By taking expectations and applying Lemma 3.2, we have
$\mathbb{E}[(X_{B})_{r}]=\sum_{d=1}^{r}\frac{N_{r,d}(B)}{p^{d}},$ (12)
where $N_{r,d}(B)$ denotes the number of ordered $r$-tuples of distinct pairs $(S_{j},b_{j})$ for which the incidence matrix has rank $d$ (over $\mathbb{F}_{p}$) and the corresponding system is consistent. By Lemma 2.1, we may freely interpret ranks over $\mathbb{Q}$.
The main term in the expression (12) of $\mathbb{E}[(X_{B})_{r}]$ comes from $d=r$. To see this, we need a crude bound on the number of low-rank incidence “patterns” as follows.
Lemma 3.3.
Fix $r\leq R$ and $d<r$. Let $T_{r,d}$ be the number of ordered $r$-tuples of distinct nonempty subsets $(S_{1},\dots,S_{r})$ of $[k]$ whose incidence matrix has rank $d$ over $\mathbb{Q}$. Then
$T_{r,d}\leq 2^{r^{2}}\Bigl(\tfrac{3}{4}\,2^{d}\Bigr)^{k}.$
Proof.
Let $V$ be an incidence matrix of rank $d<r$ and let $W=\mathrm{Col}_{\mathbb{Q}}(V)\subseteq\mathbb{Q}^{r}$. The number of possible $W$ is at most the number of ordered $d$-tuples of vectors in $\{0,1\}^{r}$ spanning $W$, hence at most $(2^{r})^{d}\leq 2^{r^{2}}$. Now fix $W$. Every column of $V$ must lie in $\Gamma:=W\cap\{0,1\}^{r}$. By Lemma 2.2 and Lemma 2.3, we have $|\Gamma|\leq\frac{3}{4}\,2^{d}$. Thus the number of possible column sequences is at most $|\Gamma|^{k}$, giving the stated bound. ∎
We now establish (11). Fix $1\leq r\leq R$. Using (12) we express $\mathbb{E}[(X_{B})_{r}]$ as a sum of contributions from the full-rank case $d=r$ and the low-rank cases $1\leq d<r$.
We first bound the low-rank cases. Let $d<r$. There are at most $m^{r}$ ways to select elements $b_{1},...,b_{r}\in B$, so we have $N_{r,d}(B)\leq m^{r}T_{r,d}$. By Lemma 3.3, we have
$\sum_{d<r}\frac{N_{r,d}(B)}{p^{d}}\leq m^{r}\,2^{r^{2}}\Bigl(\tfrac{3}{4}\Bigr)^{k}\sum_{d<r}\mathopen{}\mathclose{{\left(\frac{2^{k}}{p}}}\right)^{d}.$
Because $r\leq R\leq(\log p)^{\beta}$ with $\beta<1/2$, we have $2^{r^{2}}=p^{o(1)}$. Moreover, by Lemma 2.4, we have $2^{k}/p\leq 2\lambda\ll(\log p)^{\alpha}$, so it follows that
$\max_{d<r}\mathopen{}\mathclose{{\left(\frac{2^{k}}{p}}}\right)^{d}\leq(\log p)^{\alpha r}=\exp\bigl(O(r\log\log p)\bigr)=p^{o(1)}.$
Finally, we have $(3/4)^{k}=p^{-\eta_{0}+o(1)}$ with $\eta_{0}:=-\log_{2}(3/4)>0$. Hence, by Lemma 2.5, there exists a fixed $\eta>0$ such that uniformly for all $r\leq R$ and all $|B|=m\in\{1,2\}$,
$\sum_{d<r}\frac{N_{r,d}(B)}{p^{d}}\leq m^{r}p^{-\eta_{0}+o(1)}\leq\frac{(mM)_{r}}{p^{r}}\,p^{-\eta}.$ (13)
For the full-rank case, note that when $\operatorname{rank}(V)=r$ the map $T$ (defined in the proof of Lemma 3.2) is surjective, so the system is consistent for every right-hand side. Thus the contribution from the full-rank case equals
$\frac{1}{p^{r}}\cdot\#\{(S_{1},b_{1}),\dots,(S_{r},b_{r}):\ \text{ all distinct pairs and }\operatorname{rank}(V)=r\}.$ (14)
To estimate (14), we need to show that almost all tuples of distinct pairs $(S_{1},b_{1}),...,(S_{r},b_{r})$ satisfy $\operatorname{rank}(V)=r$. For fixed $r\leq R$ and $B\subseteq\mathbb{F}_{p}^{\times}$ with $|B|=m\in\{1,2\}$, let $\mathcal{T}_{r}(B)$ be the set of ordered $r$-tuples of distinct pairs $(S_{i},b_{i})$ with $\varnothing\neq S_{i}\subseteq[k]$ and $b_{i}\in B$.
Lemma 3.4.
Uniformly for all $1\leq r\leq R$ and $B\subseteq\mathbb{F}_{p}^{\times}$ with $|B|=m\in\{1,2\}$, the number of tuples in $\mathcal{T}_{r}(B)$ with $\mathrm{rank}(V)<r$ is at most $(mM)_{r}\,p^{-\eta_{1}}$ for some absolute constant $\eta_{1}>0$.
Proof.
We split the proof into two cases.
Case (i). Some subset repeats. This can only happen when $m=2$, and if $S_{i}=S_{j}$ then necessarily $b_{i}\neq b_{j}$ because the pairs are distinct. Fix $1\leq i<j\leq r$. There are at most $M$ choices for $S_{i}=S_{j}$, at most $2$ choices for $(b_{i},b_{j})\in B^{2}$ with $b_{i}\neq b_{j}$, and at most $(2M-2)_{r-2}\leq(2M)^{r-2}$ choices for the remaining pairs. Summing over $i<j$ gives
$\#\{\text{tuples with a repeated subset}\}\leq\binom{r}{2}\cdot M\cdot 2\cdot(2M-2)_{r-2}=\binom{r}{2}\cdot\frac{(2M)_{r}}{2M-1}\ll\frac{r^{2}}{M}\,(2M)_{r}.$
Since $r\leq R=(\log p)^{\beta}$ and $M=p^{1+o(1)}$, we have $r^{2}/M=p^{-1+o(1)}$, so this number is at most $(2M)_{r}\,p^{-1/2}$ for all large $p$.
Case (ii). All subsets are distinct. For each ordered $r$-tuple of distinct nonempty subsets $(S_{1},\dots,S_{r})$ with $\mathrm{rank}(V)=d<r$, there are at most $m^{r}$ choices of $(b_{1},\dots,b_{r})\in B^{r}$. Recall that $T_{r,d}$ is the number of ordered $r$-tuples of distinct nonempty subsets $(S_{1},\dots,S_{r})$ of $[k]$ whose incidence matrix $V$ has rank $d$. Hence
$\{\text{tuples with all }S_{i}\text{ distinct and }\mathrm{rank}(V)=d\}\ \leq\ m^{r}\,T_{r,d}.$
By Lemma 3.3,
$\sum_{d<r}m^{r}T_{r,d}\ \leq\ m^{r}\,2^{r^{2}}\Bigl(\tfrac{3}{4}\Bigr)^{k}\sum_{d<r}(2^{k})^{d}\ \leq\ m^{r}\,2^{r^{2}}\Bigl(\tfrac{3}{4}\Bigr)^{k}\cdot r\,(2^{k})^{r-1}.$
Using $M=2^{k}-1\asymp 2^{k}$ and (2), we know that $(mM)_{r}\asymp(mM)^{r}\asymp m^{r}(2^{k})^{r}$ uniformly for $r\leq R$. Let $\eta_{2}:=-\log_{2}(3/8)>0$. Putting the above bounds together, for all large $p$ we have
$\frac{\#}{(mM)_{r}}\leq\ r\,2^{r^{2}}\Bigl(\tfrac{3}{4}\Bigr)^{k}\cdot\frac{(2^{k})^{r-1}}{(2^{k})^{r}}\ =\ r\,2^{r^{2}}\Bigl(\tfrac{3}{8}\Bigr)^{k}\ =\ p^{-\eta_{2}+o(1)}\leq p^{-\eta_{2}/2}.$
Combining both cases, we conclude that the lemma holds with $\eta_{1}:=\min\{1/2,\eta_{2}/2\}$. ∎
Fix $B\subseteq\mathbb{F}_{p}^{\times}$ with $|B|=m\in\{1,2\}.$ Since $|\mathcal{T}_{r}(B)|=(mM)_{r}$, by (14) and Lemma 3.4, the contribution of the full-rank case in (12) equals
$\frac{\#\{(S_{i},b_{i})\in\mathcal{T}_{r}(B):\ \mathrm{rank}(V)=r\}}{p^{r}}=\frac{(mM)_{r}-O\mathopen{}\mathclose{{\left((mM)_{r}p^{-\eta_{1}}}}\right)}{p^{r}}=\frac{(mM)_{r}}{p^{r}}\,(1+O(p^{{-\eta_{1}}})),$
uniformly for all $1\leq r\leq R$. Together with (12) and (13) this yields
$\mathbb{E}[(X_{B})_{r}]=\frac{(mM)_{r}}{p^{r}}\bigl(1+O(p^{-\eta_{3}})\bigr)\qquad(1\leq r\leq R),$
uniformly for all $B\subseteq\mathbb{F}_{p}^{\times}$ with $|B|\in\{1,2\}$, where $\eta_{3}:=\min\{\eta/2,\eta_{1}/2\}>0$. By (2), we have $(mM)_{r}=(mM)^{r}(1+o(1))$. Since $M/p=\lambda$, we derive that
$\mathbb{E}[(X_{B})_{r}]=(m\lambda)^{r}\bigl(1+O(p^{-\eta_{3}})\bigr)\qquad(r\leq R),$ (15)
uniformly for all $B\subseteq\mathbb{F}_{p}^{\times}$ with $|B|\in\{1,2\}$, proving (11).
Now, set $\widetilde{R}:=R-1$. Applying (10) with truncation level $\widetilde{R}$ and $Z=X_{B}$ gives
$\mathbb{P}(X_{B}=0)=\sum_{r=0}^{\widetilde{R}}(-1)^{r}\frac{\mathbb{E}[(X_{B})_{r}]}{r!}+O\mathopen{}\mathclose{{\left(\frac{\mathbb{E}[(X_{B})_{\widetilde{R}+1}]}{(\widetilde{R}+1)!}}}\right).$
Since $\widetilde{R}+1=R$, we may use (15) for every $r\leq R$ to obtain, uniformly in $B$ with $|B|\in\{1,2\}$,
$\sum_{r=0}^{\widetilde{R}}(-1)^{r}\frac{\mathbb{E}[(X_{B})_{r}]}{r!}=\sum_{r=0}^{\widetilde{R}}(-1)^{r}\frac{(m\lambda)^{r}}{r!}+O\mathopen{}\mathclose{{\left(p^{-\eta_{3}}\sum_{r=0}^{\widetilde{R}}\frac{(m\lambda)^{r}}{r!}}}\right),$
and by (11) we have
$\frac{\mathbb{E}[(X_{B})_{\widetilde{R}+1}]}{(\widetilde{R}+1)!}=\frac{\mathbb{E}[(X_{B})_{R}]}{R!}=\frac{(m\lambda)^{R}}{R!}\,(1+o(1)).$
Using $\sum_{r=0}^{\widetilde{R}}(m\lambda)^{r}/r!\leq e^{m\lambda}$ and $e^{m\lambda}=p^{o(1)}$ (by Lemma 2.4), the error $O(p^{-\eta_{3}}e^{m\lambda})$ is $o(e^{-m\lambda})$. Moreover, since $\lambda=o(R)$, Stirling’s formula yields
$\frac{(m\lambda)^{R}}{R!}\leq\mathopen{}\mathclose{{\left(\frac{em\lambda}{R}}}\right)^{R}=o(e^{-m\lambda}).$
[…]
$\sum_{r=0}^{\widetilde{R}}(-1)^{r}\frac{(m\lambda)^{r}}{r!}=e^{-m\lambda}+O\mathopen{}\mathclose{{\left(\frac{(m\lambda)^{\widetilde{R}+1}}{(\widetilde{R}+1)!}}}\right)=e^{-m\lambda}+o(e^{-m\lambda}).$
Combining the above estimates, we conclude
$\mathbb{P}(X_{B}=0)=e^{-m\lambda}\,(1+o(1)),$
uniformly for all $B\subseteq\mathbb{F}_{p}^{\times}$ with $|B|=m\in\{1,2\}$. This proves Proposition 3.1.∎
4. Concluding Remarks
Our main theorem shows that, for primes $p$,
$f(p)\geq\log_{2}p+\mathopen{}\mathclose{{\left(\frac{1}{2\log 2}+o(1)}}\right)\log\log p.$ (16)
This demonstrates that the upper bound $f(N)\leq\log_{2}N+o(\log\log N)$ cannot hold uniformly over all finite abelian groups of order $N$. Nevertheless, certain abelian groups $G$ of order $N$ do exist for which, with high probability, a uniformly random subset of size $\log_{2}N+o(\log\log N)$ (or even smaller size) generates all elements of $G$. For such examples, we refer to Sothanaphan [12], which treats the case $G=(\mathbb{Z}/2\mathbb{Z})^{d}$.
Erdős and Rényi [8, Theorem 2] proved that for all $N$,
$f(N)\leq\log_{2}N+\frac{1}{\log 2}\log\log N+O(1).$ (17)
It is therefore natural to ask for an asymptotically sharp second-order term.
Problem 4.1.
Determine the following constant $c_{*}$ satisfying
$c_{*}=\limsup_{N\to\infty}\frac{f(N)-\log_{2}N}{\log\log N}$
Using (16) and (17), we can derive that $\frac{1}{2\log 2}\leq c_{*}\leq\frac{1}{\log 2}$. Heuristics and numerical experiments suggest that the constant in Problem 4.1 is very likely $c_{*}=\tfrac{1}{\log 2}$. However, the present method does not reach the regime $c_{*}>\tfrac{1}{2\log 2}$. One evident bottleneck is the proof of Lemma 3.3. It would be very interesting to push the lower bound constant beyond $\tfrac{1}{2\log 2}$.
Acknowledgements
We thank Wouter van Doorn for helpful suggestions that improved the presentation of this paper. We are grateful to Nat Sothanaphan for providing relevant materials in [12] and for an elegant rewrite [11] of the original AI-generated proof. We also thank Boris Alexeev, Ingo Althöfer, Thomas Bloom, Mehtaab Sawhney, Terence Tao, and Shengtong Zhang for helpful discussions on the Erdős Problems website forum. Finally, we are grateful to Thomas Bloom for founding and maintaining the Erdős Problems website [4]. This work is supported by National Key Research and Development Program of China 2023YFA1010201, National Natural Science Foundation of China grant 12125106, and Innovation Program for Quantum Science and Technology 2021ZD0302902.